An inductance of 1 henry and a resistance of 1 ohm are connected in series with a constant EMF of E volts. If the current is initially zero, and is equal to 10 A after 5 seconds, find E.


1
Expert's answer
2022-03-07T17:20:04-0500

"L\frac{dI}{dt}+RI=E"

"1\frac{dI}{dt}+1I=E"

"\frac{dI}{dt}+I=E"

"\frac{dI}{dt}=E-I"

"\frac{dI}{E-I}=dt"

integrate both sides;

"\int \frac{dI}{E-I}=\int dt"

"-log(E-I)|_0^i=t|_0^t"

at t=5, I=10A

Replace the above conditions in the integral;

"-log(E-10)=5"

"log\frac{E}{10}=-5"

"\frac{E}{10}=e^{-5}"

"E=e^{-5}*10"

"E=0.0674V"